# Condensing Against Online Adversaries

### v1

@@ 1.3.4 Previous results interpreted in oNOSF source context
Previous impossibility results can be interpreted in the context of extracting / condensing from uniform oNOSF sources . For instance, collective coin flipping impossibility results of [ben-or_collective_1989] imply extraction impossibility results for uniform ${\left(g,\ell,n\right)}$-oNOSF sources when $n=1$. In particular, they imply:
@@
There does not exist an $\varepsilon$-extractor for uniform ${\left(g,\ell,1\right)}$-oNOSF sources where $\varepsilon=\frac{b}{2\ell}$.
@@ 1.3.5 $\varepsilon$-collective sampling
Since collective sampling lower bounds show that for any protocol, 00-error condensing beyond rate $g/\ell$ is impossible, one can naturally ask whether condensing with small error $\varepsilon$ is possible: We call this problem $\varepsilon$-collective sampling where the goal is to output a distribution which is $\varepsilon$-close to a distribution where every output has small probability.
Interpreted this way, this is exactly what protocols arising out of our condensers for uniform oNOSF sources provide: Using 1, when each player has access to $10^{4}$ random bits, there exists a simple protocol that can handle $0.49\ell$ corrupt players such that the players can collectively sample a distribution over $m=O(\ell)$ bits which is $2^{-\Omega(\ell)}$-close to having entropy $0.99m$. As far as we are aware, such a protocol cannot be implied using any other previous protocol (a lot of previous protocols are obtained through leader election protocols which do not seem useful here since the leader has access to only constant number of bits).
@@ 1.3.6 Collective coin flipping and sampling with weak random sources
1.4 Online influence
When $n$ is a large enough constant, we showed that great condensers do exist for oNOSF sources . We now study the case of small $n$. In particular, we focus on the case of $n=1$. We call such uniform ${\left(g,\ell,1\right)}$-oNOSF sources as ${\left(g,\ell\right)}$-oNOBF sources; oNOBF stands for online non-oblivious bit-fixing sources. We ask what is the exact tradeoff between $g$, $\ell$, and $\varepsilon$ for extracting / condensing from oNOBF sources .
Towards this, we introduce a new notion of influence for Boolean functions which we call online influence.
Definition 1.7 (Online influence).
and the total online influence is
We first observe that for monotone functions, online influence equals the usual notion of influence (see Lemma 7.4 for a proof).
Next, we show a Poincaré style inequality for total online influence:
Theorem 4 (Theorem 7.5, restated).
For any $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname*{Var}(f)\leq\operatorname{\mathbf{oI}}[f]\leq\sqrt{\ell\operatorname*{Var}(f)}$.
These inequalities are tight, with tightness exhibited by PARITY and MAJORITY respectively (see Example 7.6 for more details).
We then ask the natural question that for a function $f$, what is the maximum online influence out of all $n$ bits? For the usual notion of influence, this question was resolved by the famous theorem of [kahn_influence_1988] who showed there always exists a bit with influence at least $\operatorname*{Var}(f)\cdot\Omega\left(\frac{\log\ell}{\ell}\right)$. We show that surprisingly, there exists a balanced function, namely the address function, where every bit has online influence at most $O\left(\frac{1}{\ell}\right)$ (see Lemma 7.12 for a proof). This provides a separation between the usual notion of influence and online influence. Additionally, in Lemma 7.13, we show that the address function is an extractor for uniform ${\left(\ell-1,\ell\right)}$-oNOSF sources with error $1/\ell$
Lastly, we show that using this notion of online influence, we can obtain the following extraction lower bound:
Theorem 1.8 (Informal version of Theorem 7.17).
For $\varepsilon<0.01$, there do not exist extractors for ${\left(0.98\ell,\ell\right)}$-oNOBF sources with error at most $\varepsilon$.
1.5 Extracting from local oNOSF sources
Definition 1.9 (Local oNOSF sources ).
[…]
Theorem 1.10 (Using the explicit extractor of [chattopadhyay_extractors_2022]).
@@ 2 Proof overview
Our proof overview begins by outlining our new existential results for condensers in Section 2.1 that is able to handle even constant block length. Next, we present our explicit condenser results in Section 2.2 before discussing of our low-entropy to uniform oNOSF source conversion in Section 2.3. Finally, we end with a summary of the key ideas behind our results regarding online influence in Section 2.4.
@@ 2.1 Existence of oNOSF condensers for all $\ell$ and $n$
Here we crucially used the fact that there exist seeded condensers with seed length dependence $1\cdot\log(1/\varepsilon)$. Currently, we do not have explicit constructions with this dependence. We also couldn’t have used a seeded extractor since for them, the seed length dependence is $2\cdot\log(1/\varepsilon)$. For that to work, we would need to assume $g\geq 0.76\ell$.
@@ 2.3 Converting low-entropy oNOSF sources to uniform oNOSF sources
2.4 Online influence
In this subsection, we provide a brief overview of our results regarding online influence and contrast it with the established notion of influence for Boolean functions.
@@ A Poincaré inequality and extractor lower bounds
One fundamental inequality of regular influence is that of the Poincaré inequality stating that $\operatorname*{Var}(f)\leq\operatorname{\mathbf{I}}[f]$. We prove a similar result for online influence.
@@
For any $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname*{Var}(f)\leq\operatorname{\mathbf{oI}}[f]\leq\sqrt{{\ell}\operatorname*{Var}(f)}$.
It is not hard to derive extractor lower bounds for oNOBF sources from the above result. The high level idea is to collect bits with high online influence, which is guaranteed by the above theorem (using an averaging argument) to form a coalition of coordinates that has enough online influence to bias the claimed extractor. We refer the reader to Theorem 7.17 for more details.
The proof of Theorem 2.5 is based on techniques from the Fourier analysis of Boolean functions. [^5] We give a very brief recap of necessary notions from Fourier analysis of Boolean functions in Section 7.2. The following key result that implies Theorem 2.5 in a straightforward way.
@@
The proof of the above result mostly standard Fourier analytic computation and we refer the reader to Section 7 for more details.
@@
It is not hard to see that for all $i\in[{\ell}]$, $\operatorname{\mathbf{oI}}_{i}[f]\leq\operatorname{\mathbf{I}}_{i}[f]$, with equality always holding for $i={\ell}$ as an adversarial online bit in the last index can see every good bit. Moreover, we observe that for monotone functions, the notion of online influence is equivalent to that of regular influence, demonstrating that any separation between the two notions must come from non-monotone functions.
@@ 2.5 Organization
In the remainder of our paper, we give some preliminaries in Section 3 before moving on to our core results. Section 4 details our proofs for the existence of seedless condensers for oNOSF sources for all regimes of $\ell$ and $n$, while Section 5 provides proofs for our explicit constructions of condensers. Next, Section 6 shows how to handle converting low-entropy oNOSF sources to uniform oNOSF source for a broader range of parameters. In Section 7 we introduce the notion of online influence and use it to provide an extraction lower bound for oNOBF sources . We conclude with a list of some open questions in Section 8.
@@ 3 Preliminaries
In this section we give some basic background and facts used throughout our paper. We use boldfaced font to indicate a random variable such as $\mathbf{X}$. Often we will use $\circ$ or , to indicate concatenation of blocks. So if $\mathbf{X}_{1}\sim\{0,1\}^{n}$ and $\mathbf{X}_{2}\sim\{0,1\}^{n}$, then $\mathbf{X}_{1},\mathbf{X}_{2}$ will be the concatenated random variable over $\{0,1\}^{2n}$. We will use the notation $[n]$ as shorthand for $\{1,\dots,n\}$. Also all logs in this paper will have base $2$ unless stated otherwise.
@@ 3.1 Basic probability notions
We use the notation $\mathbf{X}\approx_{\varepsilon}\mathbf{Y}$. to denote the fact that $|\mathbf{X}-\mathbf{Y}|\leq\varepsilon$.
@@
Lemma 3.2 (Min-entropy chain rule).
@@ 3.2 Condensers and extractors
Recall the definition of a seeded condenser.
@@ 4 Existence of condensers for all values of $\ell,n$
We will show that there exist condensers for uniform ${\left(g,\ell,n\right)}$-oNOSF sources for almost all settings of $\ell,n$, provided $g>0.5\ell$. Observe that a uniform ${\left(g,\ell,n\right)}$-oNOSF source is also a uniform ${\left(g\cdot s,\ell\cdot s,n/s\right)}$-oNOSF source by simply dividing up all blocks into $s$ parts. This implies that as $n$ becomes smaller (relative to $\ell$), it gets harder to condense with the hardest case being $n=1$. Our condenser will also be handle the case of $n=O(1)$ and $\ell$ arbitrarily growing:
@@ 4.2 Condenser for two uniform oNOSF sources
Let $\mathbf{X}\sim\{0,1\}^{n}$ be an arbitrary flat distribution and let $\mathsf{Cond}:\{0,1\}^{n}\to\{0,1\}^{m}$ be such that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq k$. Let $G\subset[n]$ with $|G|=n-b$ be arbitrary. Let $\mathbf{X}_{G}\sim\{0,1\}^{n-b}$ be the projection of $\mathbf{X}$ onto $G$. Let $\mathbf{X}^{\prime}\sim\{0,1\}^{n}$ be the distribution where the output bits defined by $G$ equal $\mathbf{X}_{G}$ and remaining $b$ bits are deterministic functions of the $n-b$ bits defined by $G$ under the restriction that $\operatorname{Supp}(\mathbf{X}^{\prime})\subset\operatorname{Supp}(\mathbf{X})$. Then, $H_{\infty}^{\varepsilon^{\prime}}(f(\mathbf{X}^{\prime}))\geq k-b$ where $\varepsilon^{\prime}=\varepsilon\cdot 2^{b}$.
@@ 5.1 Proving the main theorem
There exists a universal constant $C$ such that for all $n,g,\ell,\varepsilon$ where $g>\ell/2$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ satisfying: for any uniform ${\left(g,\ell\right)}$-oNOSF source $\mathbf{X}$ with $g>\ell/2$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-(3C)^{\ell}\log(2\ell n/\varepsilon)$ where $m=\frac{(n/2)-(3C)^{\ell}\log(2\ell n/\varepsilon)}{3}$.

### v2

@@ 1.4.4 Previous Results Interpreted in oNOSF source context
Previous impossibility results can be interpreted in the context of extracting / condensing from uniform oNOSF sources . For instance, collective coin flipping impossibility results of [ben-or_collective_1989] imply extraction impossibility results for uniform ${\left(g,\ell,n\right)}$-oNOSF sources when $n=1$. They imply:
@@
There does not exist an $\frac{b}{2\ell}$-extractor for uniform ${\left(g,\ell,1\right)}$-oNOSF sources.
@@ 1.4.5 $\varepsilon$-Collective Sampling
Since collective sampling lower bounds show that for any protocol, 00-error condensing beyond rate $g/\ell$ is impossible, one can naturally ask whether condensing with small error $\varepsilon$ is possible. We call this problem $\varepsilon$-collective sampling, where the goal is to output a distribution which is $\varepsilon$-close to a distribution where every output has small probability.
Interpreted this way, this is exactly what protocols arising out of our condensers for uniform oNOSF sources provide: Using 1, when each player has access to $10^{4}$ random bits, there exists a simple protocol that can handle $0.49\ell$ corrupt players such that the players can collectively sample a distribution over $m=O(\ell)$ bits which is $2^{-\Omega(\ell)}$-close to having entropy $0.99m$. As far as we are aware, such a protocol is not implied by any other previous protocol. Most previous protocols are obtained through leader election protocols, which do not seem useful here since the leader has access to only constant number of bits.
@@ 2 Proof Overview
Our proof overview begins by outlining our new existential results for condensers in Section 2.1 that is able to handle even constant block length. Next, we present our explicit condenser results in Section 2.2 before discussing of our low-entropy to uniform oNOSF source conversion in Section 2.3. We present the main ideas behind our results regarding online influence and extractor lower bounds in Section 2.4. In Section 2.5, we overview our extractor constructions for oNOBF and oNOSF sources, that is based on a general transformation from leader election protocols.
@@ 2.1 Existence of oNOSF Condensers for All $\ell$ and $n$
Here (in the inequality $0.5\ell\geq 1\cdot\log(\ell n/\varepsilon_{\mathsf{sCond}})$) we crucially used the fact that there exist seeded condensers with seed length dependence $1\cdot\log(1/\varepsilon)$. Currently, we do not have explicit constructions with this dependence. We also couldn’t have used a seeded extractor since for them, the seed length dependence is $2\cdot\log(1/\varepsilon)$. For that to work, we would need to assume $g\geq 0.76\ell$.
@@ 2.3 Converting Low-Entropy oNOSF Sources to Uniform oNOSF Sources
2.4 Online Influence and Extractor Lower Bounds
In this subsection, we provide a brief overview of our results regarding online influence and sketch how they imply extractor lower bounds against oNOBF sources . We also contrast online with the established notion of influence for Boolean functions. For any function $f:\{0,1\}^{n}\to\{0,1\}$, define the function $e(f)(x)=(-1)^{f(x)}$.
@@ A Poincaré inequality and extractor lower bounds
One fundamental inequality about regular influence is the Poincaré inequality which states that $\operatorname*{Var}(f)\leq\operatorname{\mathbf{I}}[f]$. We prove a similar result for online influence.
@@
For any $f:\{0,1\}^{\ell}\to\{0,1\}$, we have $\operatorname*{Var}(e(f))\leq\operatorname{\mathbf{oI}}[f]\leq\sqrt{{\ell}\operatorname*{Var}(e(f))}$.
It is not hard to derive extractor lower bounds for oNOBF sources from the above result. The high level idea is to collect bits with high online influence, which is guaranteed by the first inequality in the above theorem (using an averaging argument) to form a coalition of coordinates that has enough online influence to bias the claimed extractor. We refer the reader to Theorem 7.19 for more details.
The proof of Theorem 2.5 is based on techniques from the Fourier analysis of Boolean functions. [^6] We give a very brief recap of necessary notions from Fourier analysis of Boolean functions in Section 7.2. The following key result implies Theorem 2.5 in a straightforward way.
@@
The proof of the above result is mostly a standard Fourier analytic computation and we refer the reader to Section 7 for more details.
@@
It is not hard to see that $\operatorname{\mathbf{oI}}_{i}[f]\leq\operatorname{\mathbf{I}}_{i}[f]$ for all $i\in[{\ell}]$, with equality always holding for $i={\ell}$ as an adversarial online bit in the last index can see every good bit. Moreover, we observe that for monotone functions, the notion of online influence is equivalent to regular influence, so any separation between the two notions must come from non-monotone functions.
@@
2.5 Extractors via Leader Election Protocols
We sketch our main idea for constructing an extractor for oNOBF sources (Theorem 8.2). Similar ideas work more generally for extracting from oNOSF sources (Theorem 8.3). As mentioned above, we use a novel connection to leader election protocols to construct extractors. We refer the reader to Section 3.3 for a quick recap of the leader election protocols.
Suppose $\pi$ is an $(r-1)$-round leader election protocol over $\ell$ players where in each round, each player sends $1$ bit and with the guarantee that if there are at most $\delta\ell$ bad players, then a good player is chosen as leader with probability $1-\epsilon$. Suppose $\mathbf{X}$ is an ${\left(g,\ell r\right)}$-oNOBF source, where $g\geq lr-\delta\ell$. We simply partition the bits of $\mathbf{X}$ into chunks $\mathbf{X}_{1},\mathbf{X}_{2},\ldots,\mathbf{X}_{r}$, where each $X_{i}$ is on $\ell$ bits, and simulate the protocol $\pi$ by using the $j$’th bit of $\mathbf{X}_{i}$ as the message of the $j$’th player in round $i$, for all $1\leq j\leq\ell$ and $11\leq i\leq r-$ 1. At the end of this simulation suppose $j^{*}\in[\ell]$ is the chosen leader. Then we output the $j^{*}$’th bit of $\mathbf{X}_{r}$ as the output of the extractor.
Briefly, the reason that the above is a valid simulation of $\pi$ is the fact that the value of any bad bit in this online setting just depends on bits that appear before it, which is allowed in the leader election protocol (where in round $i$, the message of a bad player can be any function of the messages in the same round or previous rounds). The correctness of the extractor now follows from the fact that since the number of bad players (i.e., bad bits in $\mathbf{X}$) is at most $\delta\ell$, it follows from the guarantee of the protocol that the chosen leader $j^{*}\in[\ell]$ is a good player with probability at least $1-\epsilon$, and in this case the $j^{*}$’th bit of $\mathbf{X}_{r}$ must be uniform.
We note here that in the usual definition of leader election protocols, the requirement is to select a good leader with constant probability, which is a weaker guarantee than what we need to instantiate the above plan. It turns out that we can combine leader election protocols from prior works, in particular from [feige_noncryptographic_1999] and [alonnaorcoin], to construct protocols with the stronger guarantee we require. We refer the reader to Section 9 for more details on the construction of our leader election protocols.
@@ 2.6 Organization
In the remainder of our paper, we give some preliminaries in Section 3 before moving on to our core results. Section 4 details our proofs for the existence of seedless condensers for oNOSF sources for all regimes of $\ell$ and $n$, while Section 5 provides proofs for our explicit constructions of condensers. Next, Section 6 shows how to handle converting low-entropy oNOSF sources to uniform oNOSF source for a broader range of parameters. In Section 7 we introduce the notion of online influence and use it to provide an extraction lower bound for oNOBF sources . In Section 8, we present our explicit constructions of extractors for oNOBF and oNOSF sources using a connection to leader election protocols. In Section 9, we explicitly construct the required leader election protocols. We discuss some open questions in Section 10.
In Appendix A, we consider a natural local variant of oNOSF sources and show that it is straightforward to extract from such sources using existing extractors for small-space sources.
@@ 3 Preliminaries
In this section we give some basic background and facts used throughout our paper. We use boldfaced font to indicate a random variable such as $\mathbf{X}$. Often we will use $\circ$ or , to indicate concatenation of blocks. So if $\mathbf{X}_{1}\sim\{0,1\}^{n}$ and $\mathbf{X}_{2}\sim\{0,1\}^{n}$, then $\mathbf{X}_{1},\mathbf{X}_{2}$ will be the concatenated random variable over $\{0,1\}^{2n}$. We will use the notation $[n]$ as shorthand for $\{1,\dots,n\}$. All logs in this paper will have base $2$ unless stated otherwise.
@@ 3.1 Basic Probability Notions
We use the notation $\mathbf{X}\approx_{\varepsilon}\mathbf{Y}$ to denote the fact that $|\mathbf{X}-\mathbf{Y}|\leq\varepsilon$.
@@
Lemma 3.2 (Min-entropy chain rule, [maurer1997privacy]).
@@ 3.2 Condensers and Extractors
We recall the definition of a seeded condenser.
@@
3.3 Leader Election, Collective Coin Flipping and Sampling Protocols
We formalize the definition of protocols in the full information model. Collective coin flipping protocols, leader election protocols, and collective sampling protocols are special cases of such protocols where the output domain is $[\ell]$ and $\{0,1\}$ and $\{0,1\}^{m}$ for some $m$ respectively.
Definition 3.7 (Protocol in the full information model).
A $k$-round protocol with output domain $Y$ over $\ell$ players where each player sends $n$ random bits per round is a function
$\pi:\left(\left(\{0,1\}^{n}\right)^{\ell}\right)^{k}\to Y$
that takes in the input of each of the players during each round and outputs an element from set $Y$ which is the outcome of the protocol.
Here is how the protocol operates in the presence of a set $B\subset[\ell]$ of bad players: In round $i$, each of the players from $[\ell]\setminus B$ independently output a uniformly random element from $\{0,1\}^{n}$. Let their collective outputs be $\alpha_{i}\in\left(\{0,1\}^{n}\right)^{[\ell]\setminus B}$. Then, depending on $\alpha_{1},\dots,\alpha_{i}$, the players in $B$ together output an element of ($\left(\{0,1\}^{n}\right)^{B}$. Hence, we model the strategy of the bad players as a sequence of functions $\sigma=(\sigma_{1},\dots,\sigma_{k})$, where
$\sigma_{i}:\left(\left(\{0,1\}^{n}\right)^{[\ell]\setminus B}\right)^{i}\to\left(\{0,1\}^{n}\right)^{B},$
where $\sigma_{i}$ takes in the inputs of the good players from the first $i$ rounds and maps it to the output of the bad players for round $i$. For a fixed strategy $\sigma$, the outcome of the protocol can be modeled as follows: uniform random strings $\alpha_{1},\dots,\alpha_{k}\in\left(\{0,1\}^{n}\right)^{[\ell]\setminus B}$ are chosen, and the outcome of the protocol is
$\pi(\alpha_{1}:\sigma_{1}(\alpha_{1}),\alpha_{2}:\sigma_{2}(\alpha_{1},\alpha_{2}),\dots,\alpha_{k}:\sigma_{k}(\alpha_{1},\dots,\alpha_{k})).$
We now specialize this definition to define collective coin flipping protocols
Definition 3.8 (Collective coin flipping protocol).
A collective coin flipping protocol $\pi$ is a protocol in the full information model with output domain $Y=\{0,1\}$. Furthermore, we say $\pi$ is $(b,\gamma)$ resilient if in the presence of any set $B$ of bad players with $\left\lvert B\right\rvert\leq b$, we have that $\max_{o\in\{0,1\}}\Pr[\pi|_{B}=o]\leq 1-\gamma$.
Note that when $k=1$, the protocol $\pi$ just becomes a function over $\{0,1\}^{\ell}$; such 1-round coin flipping protocols which cannot be biased by any small set of bad players are also known as resilient functions.
We also specialize the definition of protocols to define leader election protocols:
Definition 3.9 (Leader election protocol).
A leader election protocol $\pi$ is a protocol in the full information model with output domain $Y=[\ell]$, the number of players the protocol is operating on. Furthermore, we say $\pi$ is $(b,\gamma)$ resilient if in the presence of any set $B$ of bad players with $\left\lvert B\right\rvert\leq b$, we have that $\Pr[\pi|_{B}\in B]\leq 1-\gamma$.
Remark 3.10.
The definition of resilience that we use, which is standard in the leader election and collective coin flipping literature, requires only that bad players can be elected as a leader with probability at most $1-\gamma$. Our leader election protocols satisfy (and need) the stronger measure of quality that is standard in the pseudorandomness literature: that bad players are chosen with probability at most $\varepsilon$ for small $\varepsilon$.
We lastly define collective sampling protocols:
Definition 3.11 (Collective sampling protocol).
A collective sampling protocol $\pi$ is a protocol in the full information model, typically with output domain $Y=\{0,1\}^{m}$ for some $m$ which is a function of $\ell$ and $n$. The goal of collective sampling protocols is to ensure that for every output set $S\subset\{0,1\}^{m}$ with density $\mu$, in the presence of $b$ bad players, the probability that the output lies in $S$ is at most $\varepsilon$, with the goal to make $\varepsilon$ as close to $\mu$ as possible.
@@ 4 Existence of Condensers for All Values of $\ell,n$
We will show that there exist condensers for uniform ${\left(g,\ell,n\right)}$-oNOSF sources for almost all settings of $\ell,n$, provided $g>0.5\ell$. Observe that a uniform ${\left(g,\ell,n\right)}$-oNOSF source is also a uniform ${\left(g\cdot s,\ell\cdot s,n/s\right)}$-oNOSF source by simply dividing up all blocks into $s$ parts. This implies that as $n$ becomes smaller (relative to $\ell$), it gets harder to condense with the hardest case being $n=1$. Our condenser will also be able to handle the case of $n=O(1)$ and $\ell$ arbitrarily growing:
@@ 4.2 Condenser for Two Uniform oNOSF Sources
Let $\mathbf{X}\sim\{0,1\}^{n}$ be an arbitrary flat distribution and let $\mathsf{Cond}:\{0,1\}^{n}\to\{0,1\}^{m}$ be such that $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq k$. Let $G\subset[n]$ with $|G|=n-b$ be arbitrary. Let $\mathbf{X}_{G}\sim\{0,1\}^{n-b}$ be the projection of $\mathbf{X}$ onto $G$. Let $\mathbf{X}^{\prime}\sim\{0,1\}^{n}$ be the distribution where the output bits defined by $G$ equal $\mathbf{X}_{G}$ and remaining $b$ bits are deterministic functions of the $n-b$ bits defined by $G$ under the restriction that $\operatorname{Supp}(\mathbf{X}^{\prime})\subset\operatorname{Supp}(\mathbf{X})$. Then, $H_{\infty}^{\varepsilon^{\prime}}(\mathsf{Cond}(\mathbf{X}^{\prime}))\geq k-b$ where $\varepsilon^{\prime}=\varepsilon\cdot 2^{b}$.
@@ 5.1 Proving the Main Theorem
There exists a universal constant $C$ such that for all $n,g,\ell,\varepsilon$ where $g>\ell/2$, there exists an explicit condenser $\mathsf{Cond}:(\{0,1\}^{n})^{\ell}\to\{0,1\}^{m}$ satisfying: for any uniform ${\left(g,\ell\right)}$-oNOSF source $\mathbf{X}$ with $g>\ell/2$, we have $H_{\infty}^{\varepsilon}(\mathsf{Cond}(\mathbf{X}))\geq m-(3C)^{\ell}\log(2\ell n/\varepsilon)$ where $m=\frac{1}{3}\cdot\left((n/2)-(3C)^{\ell}\log(2\ell n/\varepsilon)\right)$.